A recent Consumer Report article Jenny and Stephanie were reading reported that the mean number of gallons of gas consumed by a "typical American household" in a year is 750 and the standard deviation is 37 gallons. Assuming the data are normally distributed, what percentage of American households consume less that 713 gallons?
Answer by stanbon(75887)

A recent Consumer Report article Jenny and Stephanie were reading reported that the mean number of gallons of gas consumed by a "typical American household" in a year is 750 and the standard deviation is 37 gallons. Assuming the data are normally distributed, what percentage of American households consume less than 713 gallons?
----------
z(713) = (713-750)/37 = -1
------
P(x < 713) = P(z < -1) = normalcdf(-100,-1) = 0.1587 = 15.87%
================================
Cheers,
Stan H.
==============